compute the value for the given investment using the different types of compounding. present value is $8,000. Annual interest rate is 4.5% for 3 years. Compounded: a. Semi Annually b. quarterly c. monthly and d. daily
Answer by jorel1380(3719)

8000 @ 4.5% for 3 years=
8000x(1.045)^3=8000x1.141166125=$9129.33
a. Semi-annually=
8000 x (1+(0.045/2))^6=8000 x (1.0225)^6=
8000 x 1.142825441564707275390625=$9142.60
b. Quarterly=
8000 x (1+(.045/4))^12=8000 x 1.01125^12=8000 x 1.14367444074=
$9149.3955
c. Monthly=
8000 x (1+(.045/12)^36=8000 x 1.0035^36=8000 x 1.1442478322=
$9153.98266
d. Daily=
8000 x (1.00012328767)^1095=8000 x 1.14452726043=
$9156.21808346